Sue is now seven times as old as louise. in six years, sue will be five times as old as louise. find the present age of each one
Answer by checkley77(12844)

Louise=x
Sue=7x
7x+6=5(x+6)
7x+6=5x+30
7x-5x=30-6
2x=24
x=24/2
x=12 age of Louise now.
7*12=84 age of Sue now.
Proof:
7*12+6=5(12+6)
84+6=5*18
90=90